OpenStudy (anonymous):

2. Jacob spends 60 minutes in the gym every day doing freehand exercises and running on the treadmill. He spends 30 minutes more running on the treadmill than doing freehand exercises. Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does free hand exercises(y) and the number of minutes he runs on the treadmill(x). (5 points) Part B: How much time does Jacob spend on doing freehand exercises? (3 points) Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. @mathmath333

OpenStudy (anonymous):

Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does free hand exercises(y) and the number of minutes he runs on the treadmill(x). (5 points) Part B: How much time does Jacob spend on doing freehand exercises? (3 points) Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points)

OpenStudy (anonymous):

Can you help me @mathmath333 @magepker728

OpenStudy (mathmath333):

jacob does two things in gym freehand exercises and running on the treadmill total time he spends is \(60~ min\) if the time he spends for freehand exercises =\(x\) then the time he spends for running on the treadmill.=\(x+30\) hence \(\large\tt \begin{align} \color{black}{ x+ (x+30)=60\hspace{1.5em}\\ }\end{align}\)

OpenStudy (anonymous):

is this for part A

OpenStudy (anonymous):

????

OpenStudy (mathmath333):

it is for all A B and C indirectly

OpenStudy (anonymous):

Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points) can x= 40 ? we have x=y+30 or 40= y+30 --> y= 10 minutes doing freehand x+y= 40+10 = 50 minutes He stays at the gym for 60 minutes. It is possible he goofs off for 10 minutes, and does exercises for 50 minutes. So I would say yes, he could have done 40 minutes of running.

OpenStudy (mathmath333):

haha

OpenStudy (anonymous):

can you help me with b

OpenStudy (mathmath333):

did u find \(x\)

OpenStudy (anonymous):

does x=30

OpenStudy (mathmath333):

no

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it equals 15

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ x+ (x+30)=60\hspace{1.5em}\\~\\ 2x+30=60\hspace{1.5em}\\~\\ 2x=60-30\hspace{1.5em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

yes thats B

OpenStudy (anonymous):

what about c

OpenStudy (anonymous):

Part B: Plug in your value for x remember: x=y+30 so (y+30)+y=60 y+30+y=60 2y+30=60 2y=30 y=15 so 15 minutes doing freehand exercises so your right...c would be.....hold on

OpenStudy (mathmath333):

first what did find for \(A\)

OpenStudy (anonymous):

y=15 x=45

OpenStudy (anonymous):

Jacob does not spend 30 minutes on the treadmill, he spends 30 minutes MORE than 15, which is 45. 15+45=60 y+x=60

OpenStudy (anonymous):

I think I'm right not sure though...

OpenStudy (anonymous):

yeah I think that's right :D LOL

OpenStudy (mathmath333):

if \(\large\tt \begin{align} \color{black}{(x+30)=40\hspace{1.5em}\\~\\ x=40-30\hspace{1.5em}\\~\\ x=40-30\hspace{1.5em}\\~\\x=10 }\end{align}\) as \(x\) is positive integers so \(\Large\tt yes \) he can possibly spent 40 minutes running on the treadmill

OpenStudy (anonymous):

thank you guys sooooooooooo much

OpenStudy (anonymous):

welcome:)